### v1

# Channel Knowledge Map-assisted Dual-domain Tracking and Predictive Beamforming for High-Mobility Wireless Networks

Abstract
This paper introduces a novel channel knowledge map (CKM)-assisted dual-domain tracking and predictive beamforming scheme for high-mobility wireless networks. The central premise is that the CKM integrates both the coordinate and beam domains, thereby enabling tracking in one domain via treating the other domain’s input as priors or measurements. In the coordinate domain (C-Domain), an extended Kalman filter (EKF) is employed to predict and track the state (i.e., location and velocity) of a moving communication receiver across time slots under both line-of-sight (LoS)-present and LoS-absent conditions, where the CKM provides a prior mapping from multipath channel parameters to potential target locations. In the beam domain (B-Domain), the updated location of the receiver is fed back to CKM to offer a priori information of angle of arrival (AoA) variations, which are incorporated to establish beam transition models for effective beam tracking, depending on the angular variation situation of each path. Then, we analyze the Cramér-Rao Bound (CRB) for AoA estimation for each path in the considered system and propose a jointly predictive beamforming and power allocation design to minimize AoA estimation errors, directly enhancing multipath beam tracking accuracy and indirectly improving target tracking performance. Simulation results demonstrate that the proposed scheme achieves significant improvements in both target and beam tracking performance compared to the state-of-the-art approaches, particularly in AoA tracking of non-line-of-sight (NLoS) paths, highlighting the potential gain of CKM in facilitating both target and beam tracking in high-mobility communications.
Keywords:
Tracking, dual-domain, channel knowledge map, extended Kalman filter, high-mobility, predictive beamforming.
I Introduction
I-A Background
The upcoming sixth-generation (6G) wireless networks are expected to achieve unprecedented data rates, increased connection density, and superior reliability [2]. A critical aspect of 6G evolution lies in supporting high-mobility scenarios, which are vital for emerging applications involving unmanned aerial vehicles (UAVs), ground vehicles, and high-speed trains [3, 4]. In practice, these high-mobility scenarios introduce formidable challenges for accurate channel estimation or prediction due to the rapid time-varying channel conditions, further complicated by the expansion of the channel matrix across both frequency and spatial domains. The task of estimating these dynamic and large-scale channel matrices is computationally-intensive and requires a large amount of channel training overhead. Therefore, various advanced techniques such as beam tracking and predictive beamforming (BF) are essential to maintain high-quality links in high-mobility wireless networks, enabling real-time system adaptation to dynamic channel variations [5, 6, 4, 7]. Moreover, integrated sensing and communication (ISAC) has been recognized as one of the six key application scenarios for 6G [8]. In fact, leveraging communication signals for target tracking in high-mobility wireless networks not only meets emerging demands for localization services but also improves communication performance, particularly when the moving target is also the communication terminal.
One of the key challenges in high-mobility wireless networks is tracking the angles of arrival (AoA) of moving communication terminals, such as vehicles in vehicle-to-everything (V2X) networks. In particular, multipath AoA information is essential for both channel reconstruction and wireless positioning. However, AoA tracking in multipath environments is much more complicated than that in line-of-sight (LoS) propagation environments, particularly for non-line-of-sight (NLoS) paths. Specifically, NLoS paths typically experience higher propagation loss than that of the LoS path, leading to reduced receiving signal strength and AoA estimation accuracy. Moreover, the rapid changes in the AoA, driven by vehicle maneuvers and the occurrence of random obstacles, pose significant challenges for tracking multipath AoA. In addition, the fast AoA variation leads to a severe model mismatch between the assumed beam transition model and the ground truth. Even worse, the random path birth and death, or burst variations of multipath AoA significantly reduce the temporal correlation of beam transition and thus might cause the traditional beam tracking algorithm to fail [9]. As a result, incorporating extra environment information is essential to ensure reliable and accurate beam tracking in multipath environments, yet it has not been fully exploited.
Target tracking has long been a cornerstone of radar technology research [6, 10], and has recently regained significant interest in the wireless communication sector due to the emerging of the ISAC paradigm. Traditionally, radar-based target tracking usually assumes LoS propagation and ignores multipath effects, as radar typically operates in open-air spaces with extensive sensing ranges. However, ISAC operates on wireless infrastructures in complicated propagation environments, such as urban, where the sensing range is usually limited within a cell, say 500 meters [11]. Therefore, multipath effects and the random absence of the LoS path are inevitable for target tracking in the context of ISAC. Unfortunately, when the LoS path is obscured, the NLoS paths, caused by scatterers in the propagation environment, are not directly related to the moving target, thus complicating their explicit exploitation for target tracking. In this case, acquiring comprehensive environmental information becomes essential for effective beam tracking and target tracking in NLoS propagation environments.
Channel knowledge map (CKM) is a novel technique connecting environmental information and channel state information (CSI) for a specific area [2, 12, 13]. Beyond facilitating an understudy of channel temporal/frequency correlations, the CKM provides more channel a priori information by mapping geographic positions to CSI, which enables more accurate and robust beam tracking performance [14], especially in NLoS propagation environments. However, the challenge remains that location information for a moving terminal is usually unavailable at the transmitter, and the method to effectively integrate channel temporal correlations and the a priori channel information provided by the CKM remains unclear. Furthermore, the CKM can also be inversely exploited to improve target localization and tracking performance by mapping the estimated CSI to geographic a priori information. In particular, for the case with the absence of LoS path, CKM can provide geographic a priori information via exploiting the estimated NLoS path parameters and thus might improve sensing performance in NLoS propagation environment. However, obtaining accurate multipath information in high-mobility wireless networks is itself a challenging task. Therefore, this paper aims to propose a dual-domain tracking scheme, which simultaneously performs target tracking in the coordinate domain (C-Domain) and beam tracking in the beam domain (B-Domain), while bridging these two domains through the CKM.
In high-mobility wireless networks, predictive BF is critical to maintaining consistent link quality, as the procedures for CSI acquisition and BF design require substantial training overhead. In particular, based on predicted CSI or multipath AoA information, the transmitter designs BF strategies preemptively that optimize system communication and/or sensing performance for upcoming time slots. However, most existing works on predictive BF algorithms assume the presence of a LoS path [15], a condition that often dues not hold in complex environments, such as urban vehicular networks. In these settings, LoS path might be obstructed by tall buildings or other vehicles and NLoS paths are typicality inevitable. Moreover, as analyzed in [7], BF directly towards the AoA of the LoS path is not always the optimal strategy in terms of achievable rate, with a performance gap of nearly one-third compared to optimal BF in a multipath channel. Indeed, effective predictive BF design in NLoS propagation environments enables channel utilization and provides a stronger received signal/echo power, which in turn improves the beam tracking performance. Building on this insight, this paper further designs a predictive BF strategy that provides high-quality measurements for the proposed CKM-assisted dual-domain tracking scheme.
I-B Existing Works
Beam tracking has been extensively studied in literature for high-mobility wireless networks, though most existing studies have relied on LoS-dominant channel model. To mitigate the substantial overhead, which can reach up to $43.24\%$ in fifth-generation (5G) New Radio (NR) networks [5], the echo of the downlink signal reflected from the moving terminal is directly processed to estimate its state parameters (e.g., range, velocity, AoA, etc.). For instance, in [6], the authors proposed a beam tracking algorithm leveraging the temporal correlation of AoA, modeling the AoA variation within a single transmission frame as a discrete Markov process. Furthermore, in [16], an “on-grid” AoA transition model was adopted for simplicity, assuming that the AoA of each path shifts among some predetermined angular grid points. Additionally, in [17], the beam tracking problem was formulated as a Markov decision process and the corresponding training beam sequence was optimized by exploiting a reinforcement learning (RL) framework, albeit at a high computational cost. However, the adopted narrow training beams [17] are inadequate for tracking rapid beam variations in high-mobility scenarios. To overcome this, multiple radio frequency (RF) chains were advocated to generate multiple probing beams for improving tracking performance. On the other hand, for target tracking, in [18], an extended Kalman filter (EKF) framework was proposed to track the location and angular variation of a vehicle, based on echo signal measurements and the vehicle’s state evolution model. In [19], the authors proposed a Bernoulli Gaussian sum filter for tracking dynamic targets that randomly appear and disappear, especially in multi-sensor scenarios. Additionally, the unscented Kalman filter (UKF) [20] and the particle filter [21] have also been proposed for tracking moving targets in high-mobility wireless networks.
In practice, the LoS path is not always available and NLoS paths are more frequently encountered in wireless channels, which imposes challenges for both beam and target tracking, as previously discussed. Although few works have studied the beam and target tracking problems in NLoS propagation environments [9, 4, 22, 23, 24], research remains limited. In [9, 4], CKM was adopted to assist LoS path detection, which improves the robustness of target tracking via disabling state update when the LoS path is absent. Also, in [22], the authors leveraged the probability density function of the state of a moving target to mitigate NLoS propagation effects, employing an EKF to track the target from identified LoS time-of-arrival (TOA) measurements. Furthermore, in [23], the authors proposed a target tracking scheme for NLoS environments by jointly estimating the target trajectory and sensor-target distance via multi-sensor information fusion. Moreover, in [24], deep learning (DL) was applied to forecast AoAs in NLoS channels, which requires a large amount of data and incurs an expensive computation load due to the need for neural network training. The primary disadvantages of existing beam tracking methods in NLoS propagation environments include poor performance due to high propagation loss and the impracticality of the commonly-used NLoS model, which neglects the dynamic appearance and disappearance of NLoS paths caused by random obstacles.
Predictive BF has emerged as a key strategy to mitigate frequent link blockages and reduce significant beam training overhead in high-mobility communication systems. In practice, moving communication terminals in such wireless networks often traverse fixed routes, such as high-speed rails or expressways, providing predictable position and velocity, thus making predictive BF feasible. For example, in [25], the authors introduced a low-complexity predictive BF scheme featuring transmit diversity, while the BF vector associated with the predictive location of a moving communication receiver is precalculated and pre-stored at the base station. In contrast, [26] introduced a vision-aided approach leveraging camera images and DL to predict the optimal beam indices in the next few time slots. Moreover, in [27], the authors trained DL models based on historical CSI to forecast future beam patterns. However, both approaches in [26] and [27] require a large amount of data and a high computational demand for training DL models.
Wireless CKM-assisted sensing expands the design possibilities by providing an interface between beam and coordinate domains. In [28], the authors proposed a LoS-map-assisted anchor selection scheme to minimize the Bayesian Cramér-Rao Lower Bound (BCRLB) for ranging. In addition, the authors in [29] proposed a clutter angle map (CLAM) to eliminate the clutter-related components from the sensing signal before target detection, thereby enhancing target sensing performance. To the authors’ best knowledge, CKM has not been fully exploited to assist effective beam tracking in multipath channels but has been exploited in the opposite way to improve target tracking performance. Indeed, CKM connects the C-Domain and B-Domain, potentially improving tracking performance in both domains by treating the information from one domain as measurements for the other. This integration motivates the proposed dual-domain tracking scheme in this paper.
I-C Our Contributions
To tackle the challenges posed by multipath propagation and high-mobility networks, this paper proposes a CKM-assisted dual-domain tracking framework. The proposed scheme simultaneously performs target tracking in the C-Domain and beam tracking in the B-Domain by modeling the state evolution in one domain and utilizing measurements/a priori information from the other domain. To further improve beam tracking performance, this paper proposes a predictive BF design to minimize the maximum AoA estimation errors across multiple paths. Our main contributions are summarized as follows:
  • We propose a CKM-assisted dual-domain tracking framework that effectively integrates information from both the coordinate and beam domains. In the C-Domain, the location and velocity of the target are the states to be updated, with the estimated channel parameters from the B-Domain acting as measurements, directly for LoS-present cases and indirectly via CKM for LoS-absent cases. In the B-Domain, the received echo serves as the measurement and the multipath AoAs are the state to be updated, via combining the a priori information provided by the C-Domain through CKM. This integration can enhance beam tracking accuracy by exploiting CKM to reference the tracked state of the moving target and improve target tracking accuracy by inversely leveraging CKM with the tracked CSI.
  • In the C-Domain, we propose an EKF-based target tracking algorithm to predict and track the state of a moving user terminal, with state evolution and measurement models defined for both LoS-present and LoS-absent conditions. Even when the LoS path is obscured, the proposed approach can maintain continuous target tracking by exploiting NLoS path information through the CKM. In the B-Domain, depending on the condition of the beam transition, we propose an adaptive a priori information fusion strategy for creating the beam transfer probability matrices (TPMs) to assist beam tracking, which effectively combines the channel temporal correlation and the channel a priori information provided by CKM. The proposed approach can track the fast and even burst changes of multipath AoA due to the dynamic environment by exploiting CKM.
  • Building upon the predicted AoA, we analyze the Cramér-Rao Bound (CRB) of AoA estimation for each path to accurately characterize the AoA estimation error. Then, a predictive BF and power allocation design is formulated as an optimization problem to minimize the maximum AoA estimation error among multiple paths for the next time slot. We propose a suboptimal approach that first selects beams based on the predicted AoAs and then designs the power allocation among multiple beams via solving a convex optimization problem.
  • We conduct extensive simulations to evaluate the superior performance of the proposed dual-domain tracking and predictive BF scheme. Our simulation results demonstrate that the proposed scheme achieves both accurate and continuous target and beam tacking even in complex high-mobility wireless networks, while traditional tracking schemes often fail. Besides, the proposed predictive BF and power allocation design further improves the beam tracking accuracy, particularly for NLoS paths.
The remainder of this article is organized as follows: Section II introduces the system model. Section III presents the proposed CKM-assisted dual-domain tracking framework. Sections IV and V describe the proposed C-Domain and B-Domain tracking algorithms, respectively. Section VI presents a predictive BF design. Section VII presents the numerical results and Section VIII concludes the paper.
Unless otherwise specified, matrices are denoted by uppercase bold letters and vectors are represented by lowercase bold letters; $\operatorname{vec}(\cdot)$ denotes the vectorization operations; $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{*}$ stand for transpose, Hermitian transpose, and the complex conjugate of a matrix; $\boldsymbol{I}_{N}$ denotes the $N$-dimensional identity matrix; $\boldsymbol{1}_{N}$ denotes the $N$-dimensional all-ones vector; $\delta(\cdot)$ denotes the Dirac delta function; $\mathbb{C}$ denotes a complex space; $\mathbb{E}$ denotes the expectation operator; $\mathrm{Var}(\cdot)$ and $\mathrm{Cov}(\cdot)$ denote the operator to determine the variance and covariance, respectively; $\operatorname{diag}(\cdot)$ denotes the diagonal operator; $\hat{x}$ denotes the estimation of variable $x$; $\otimes$ denotes the Kronecker product; $\propto$ indicates that the left hand side is proportional to the right hand side; $\boldsymbol{x}\sim\mathcal{CN}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes a circularly symmetric complex Gaussian vector with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$; $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes a real-valued Gaussian vector with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$; $\boldsymbol{A}_{k,:}$, $\boldsymbol{A}_{:,l}$ and $\boldsymbol{A}_{k,l}$ represent the $k$-th row, $l$-th column and $(k,l)$-th element of $\boldsymbol{A}$; $l_{1}$ norm, $l_{2}$ norm, and the Frobenius norm are denoted by $\|\cdot\|_{1}$, $\|\cdot\|_{2}$, and $\|\cdot\|_{F}$, respectively.
II System Model
Consider a downlink high-mobility wireless vehicular network [^1] Note that the considered vehicular network is a typical high-mobility network and the proposed dual-domain tracking framework can be extended to other high-mobility wireless networks, such as UAV networks. as depicted in Fig. 1. A roadside unit (RSU) serves as both a communication transmitter and a sensing receiver, providing services to a moving vehicle. The full-duplex operation RSU is equipped with $N_{t}$ transmitting antennas and $N_{r}$ receiving antennas, both constituting uniform linear arrays (ULAs) with half wavelength antenna spacing. The vehicle is equipped with a ULA of $N_{s}$ antenna elements for receiving signals from the RSU. For ease of exposition, as adopted in the literature [18, 9, 5, 7], we assume that the vehicle travels along the road parallel to the antenna array of RSU, initially opposite and subsequently moving away from the RSU [^2] To facilitate the presentation of the proposed dual-domain tracking framework, we consider a simple tracking scenario in this paper. However, it is worth noting that the proposed framework can be extended to accommodate more complicated tracking scenarios, such as turning vehicles.. In the considered system model, the position of the vehicle is represented by $\boldsymbol{q}=[q_{x},q_{y}]$ and its velocity is denoted as $v_{1}$. We assume that the velocity is a scalar, with its sign indicating movement direction along the x-axis, defined as the horizontal axis in Fig. 1.
Figure 1: Beam tracking system model for a high-mobility vehicle network in multipath environments.
Assume $T_{\max}$ being the maximum time duration of interest. The time period $T_{\max}$ is divided into several time slots of length $\Delta T$, with each slot further divided into $L$ symbol periods. Within each slot, we assume that the channel parameters remain unchanged [18, 5, 30]. The transmitted $N_{s}$ data streams from the RSU in a slot are given by $\boldsymbol{S}=[\boldsymbol{s}^{T}_{1};\cdots;\boldsymbol{s}^{T}_{N_{s}}]=[\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{L}]\in\mathbb{C}^{N_{s}\times L}$, where $\boldsymbol{s}_{n_{s}}\in\mathbb{C}^{L\times 1}$, ${n_{s}}\in\{1,\ldots,N_{s}\}$, denotes the $n_{s}$-th downlink data streams and $\boldsymbol{x}_{l}\in\mathbb{C}^{N_{s}\times 1}$, denotes the transmitted signal at the $l$-th snapshot, $\forall l\in\{1,\ldots,L\}$. We assume $\boldsymbol{S}_{i,j}\sim\mathcal{CN}(0,1)$, $\forall i,j$, are independent and identically distributed (i.i.d.), leading to $\mathbb{E}[\boldsymbol{x}_{i}\boldsymbol{x}_{j}^{H}]=\delta(i-j)\boldsymbol{I}_{N_{s}}$, $\forall i,j$. Owing to the independence among different data streams, for sufficiently large $L$ we have $\boldsymbol{s}^{H}_{n_{s}}\boldsymbol{s}_{n_{s}^{\prime}}\approx\delta(n_{s}-n_{s}^{\prime})L$. The $N_{s}$ downlink data streams transmitted at the $[\eta]$-th slot are denoted as $\boldsymbol{S}[\eta]=[\boldsymbol{x}_{1}[\eta],\cdots,\boldsymbol{x}_{L}[\eta]]\in\mathbb{C}^{N_{s}\times L}$. The precoded signal in the $[\eta]$-th slot at the RSU is $\tilde{\boldsymbol{S}}[\eta]=\boldsymbol{F}[\eta|\eta-1]\boldsymbol{S}[\eta]\in\mathbb{C}^{N_{t}\times L}$, where $\boldsymbol{F}[\eta|\eta-1]\in\mathbb{C}^{N_{t}\times N_{s}}$ is the predictive BF matrix. In high-mobility scenario, obtaining perfect CSI at the transmitter is challenging. Thus, the predictive BF matrix $\boldsymbol{F}[\eta|\eta-1]$ is designed based on the predicted CSI in the proposed dual-domain tracking framework.
Assuming there are a finite number of scatterers between the transmitter and receiver, the received echoes at the RSU consist of the signals reflected by both scatterers and the moving vehicle. The sensing channel in the $\eta$-th slot is given by
$\small\boldsymbol{H}[\eta](t,\tau)=\sum_{i=1}^{P}\beta_{i}[\eta]\boldsymbol{b}(\theta_{i}[\eta])\boldsymbol{a}^{H}(\theta_{i}[\eta])\delta(\tau-\tau_{i}[\eta])e^{j2\pi\mu_{i}[\eta]t},$ (1)
where $i\in\{1,\dots,P\}$ is the path index [^3] For the LoS-present case, path 1 is a LoS path, while path 2 is an NLoS path. For LoS-absent case, path 1 becomes the strongest NLoS path., $\beta_{i}[\eta]\in\mathbb{C}$, $\tau_{i}[\eta]\in\mathbb{R}$ and $\mu_{i}[\eta]\in\mathbb{R}$ denote the channel path gain, the round-trip delay, and the round-trip Doppler frequency shift of the $i$-th path in the $[\eta]$-th slot, respectively. $\boldsymbol{a}(\theta[\eta])$ and $\boldsymbol{b}(\theta[\eta])$ denote the steering vectors of the RSU transmitting and receiving ULAs, respectively, and they are defined by $\boldsymbol{a}(\theta[\eta])=[1,e^{j\pi\cos(\theta[\eta])},\cdots,e^{j\pi(N_{t}-1)\cos(\theta[\eta])}]^{T}\in\mathbb{C}^{N_{t}\times 1}$, and $\boldsymbol{b}(\theta[\eta])=[1,e^{j\pi\cos(\theta[\eta])},\cdots,e^{j\pi(N_{r}-1)\cos(\theta[\eta])}]^{T}\in\mathbb{C}^{N_{r}\times 1}$. The channel path gains $\beta_{i}[\eta]$ captures both effects of the radar cross-section (RCS) of the vehicle/scatterers and the propagation loss. Denote $\theta_{i}[\eta]$ as the angle of the $i$-th path relative to the RSU and we assume that the AoA and angle of departure (AoD) of the $i$-th path at the RSU in the $[\eta]$-th slot are identical, as commonly adopted in [30, 18].
In the remaining part of this paper, the slot index is omitted without causing ambiguity. Accordingly, the echo signal, $\boldsymbol{R}$, received over the $L$ symbol periods at the RSU is given by:
$\small\boldsymbol{R}=\sum_{i=1}^{P}\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}{\boldsymbol{S}_{l_{i}}}{\boldsymbol{\Lambda}_{k_{i}}}+\boldsymbol{Z}\in\mathbb{C}^{N_{r}\times[L+\max(l_{i})]},$ (2)
where ${\boldsymbol{S}_{l_{i}}}\triangleq[\boldsymbol{0}_{N_{s}\times l_{i}},\boldsymbol{S},\boldsymbol{0}_{N_{s}\times[\max(l_{i})-l_{i}]}]\in\mathbb{C}^{N_{s}\times(L+\max(l_{i}))}$ denotes the delayed signal associated with the $i$-th path, $l_{i}=\frac{\tau_{i}}{T_{p}}$ represents normalized delay index within a slot and $T_{p}=\frac{\Delta T}{L}$ denotes the time domain sample interval. We assume that the delays are integer multiples of the sampling interval and there is no inter-symbol interference among consecutive time slots by inserting a proper temporal guard interval at the beginning of each time slot [18]. ${\boldsymbol{\Lambda}_{k_{i}}}$ is the Doppler frequency shift matrix and it is defined by ${\boldsymbol{\Lambda}_{k_{i}}}\triangleq\operatorname{diag}(\boldsymbol{0}_{l_{i}},\boldsymbol{\Lambda}^{\text{shift}}_{k_{i}},\boldsymbol{0}_{\max(l_{i})-l_{i}})\in\mathbb{C}^{(L+\max(l_{i}))\times(L+\max(l_{i}))}$, where $k_{i}=\mu_{i}T_{p}$ represents the normalized Doppler shift and $\boldsymbol{\Lambda}^{\text{shift}}_{k_{i}}=\operatorname{diag}(1,\dots,e^{j2\pi k_{i}(L-1)})\in\mathbb{C}^{L\times L}$. The noise signal is denoted as $\boldsymbol{Z}$ where each entry follows i.i.d. Gaussian with zero mean and variance $\sigma_{z}^{2}$, i.e., $\boldsymbol{Z}_{n_{r},:}\sim\mathcal{CN}(\boldsymbol{0},\sigma_{z}^{2}\boldsymbol{I}_{L+\max(l_{i})}),\forall n_{r}=1,\ldots,N_{r}$, and $\sigma_{z}^{2}$ is the noise power.
CKM aims to provide location-specific channel parameter tuples for an area of interest. Specifically, CKM is essentially a mapping function from the user terminal location $\boldsymbol{q}$ to the location-specific channel knowledge such as path gain, angles, delay, and Doppler shift444 In the CKM literature [12], the channel Doppler shift can be obtained from a CKM based on the location and velocity information. In practice, the Doppler information can be sampled in a dynamic environment or inferred based on the sampled Doppler shift in a static scattering environment. This paper focused on tracking algorithm design and thus the CKM construction is out of scope. Note that the proposed dual-domain tracking scheme can be straightforwardly extended to the case without Doppler information in the CKM.[12], i.e.,
$\mathcal{M}(\boldsymbol{q},v_{1})=\left\{(\alpha_{i},{\theta}_{i},\tau_{i},\mu_{i})|i=1,\cdots,P\right\}.$ (3)
where $\alpha_{i}\in\mathbb{R}$ captures the propagation power loss of the $i$-th path relative to the RSU. The channel path gain is given by $\beta_{i}=\varepsilon\alpha_{i}^{2}$, and $\varepsilon$ is the time-invariant reflection coefficient of the considered user terminal [6, 18, 4]. To facilitate the presentation in sequel, we assume that the number of paths stored in CKM is the same as the number of paths $P$ in (1). In practice, the number of paths during the time duration of interest might be time-varying and be inconsistent with the number of paths stored in the CKM, requiring further investigation, which is left for future work.
III CKM-assisted Dual-domain Tracking Framework
Before delving into the details, in this section, we present a high-level overview to the proposed CKM-assisted dual-domain tracking framework, as illustrated in Fig. 2.
Figure 2: CKM-assisted dual-domain tracking framework.
The proposed framework is well-designed with the traditional Kalman filtering framework [18, 3, 31], which consists of state prediction and update steps across consecutive time slots. Distinctively, unlike existing ones, it is defined in both the beam and coordinate domains, interconnected through the CKM.
In the C-Domain, the measurement models are addressed separately for two scenarios: LoS-present and LoS-absent conditions. Adopt the index $r=1,2$ to denote the LoS-present and LoS-absent cases, respectively. For LoS-present cases, the C-Domain measurement, vector $\boldsymbol{\beta}_{1}[\eta]=[\tau_{1}[\eta],\nu_{1}[\eta],\cos(\theta_{1}[\eta])]^{T}\in\mathbb{C}^{3\times 1}$, comprises the delay, Doppler shift, and the cosine of the AoA for the LoS path. For LoS-absent cases, vector $\boldsymbol{\beta}_{2}[\eta]=[\boldsymbol{\tau}[\eta];\boldsymbol{\mu}[\eta];\cos(\boldsymbol{\theta}[\eta])]\in\mathbb{C}^{3P\times 1}$ is defined by the B-Domain estimated path parameters ${\boldsymbol{\tau}}[\eta]$, ${\boldsymbol{\nu}}[\eta]$, and ${\boldsymbol{\theta}}[\eta]$ for all $P$ paths. On the other hand, the C-Domain states are defined identically for both LoS-present and LoS-absent cases, which consists of the position and velocity of the moving vehicle, i.e., $\boldsymbol{\alpha}[\eta]=\left[\boldsymbol{q}[\eta],v_{1}[\eta]\right]^{T}\in\mathbb{R}^{3\times 1}$. A measurement model exploiting the geometric relationship that connects $\boldsymbol{\alpha}[\eta]$ and $\boldsymbol{\beta}_{1}[\eta]$ is employed in the presence of the LoS path; otherwise, the CKM that connecting multipath parameter estimates and the state of a moving vehicle is utilized to construct the measurement model, as will be detailed in Section IV of this paper. Moreover, an identical C-Domain state evolution model, i.e., a constant velocity movement model, is adopted for both LoS-present and LoS-absent cases. In the $[\eta]$-th time slot, by leveraging the predicted vehicle state $\boldsymbol{\hat{\alpha}}[\eta|\eta-1]$ and the C-Domain measurements $\boldsymbol{\beta}_{r}[\eta],r=(1,2)$, the RSU applies an EKF-based method to update the state of the vehicle, as will be detailed in Section IV of this paper. Then, based on the C-Domain state evolution model, the state of the vehicle is predicted for the $[\eta+1]$-th slot, i.e., $\boldsymbol{\hat{\alpha}}[\eta+1|\eta]$, which is further exploited for state update in the $[\eta+1]$-th time slot.
In the B-Domain, measurement is the received echo signal at the RSU and state corresponds to the multipath AoA. At the beginning of time slot $[\eta]$, the RSU first receives echo signals $\boldsymbol{R}[\eta]$ based on the predictive BF matrix $\boldsymbol{F}[\eta|\eta-1]$, facilitating the extraction of the round trip delay $\boldsymbol{\tau}[\eta]\in\mathbb{R}^{P\times 1}$ and Doppler shift $\boldsymbol{\mu}[\eta]\in\mathbb{R}^{P\times 1}$ through the matched filtering technique [32, 33]. Note that different propagation paths possess different $\tau_{i}[\eta]$ and $\nu_{i}[\eta]$, and thereby the RSU is capable of separating the signals $\boldsymbol{R}_{i}[\eta]$ reflected from different scatterers. For the $i$-th path, the state $\{\theta_{i}[\eta]\}_{i=1}^{P}$ in the B-Domain can be updated by incorporating the state prediction $\{\boldsymbol{p}(\theta_{i}[\eta|\eta-1])\}_{i=1}^{P}$ from the previous slot and the measurement $\boldsymbol{R}_{i}[\eta]$ in the current slot, as detailed in Section V of this paper. The B-Domain state evolution model, which is characterized by the beam TPM, needs to be established from beam tracking. This model relies on the channel’s temporal correlation as well as the a prior information stored in the CKM, where the latter one further depends on the predicted state of the vehicle $\hat{\boldsymbol{\alpha}}[\eta+1|\eta]$. Specifically, in Fig. 2, the matrix $\boldsymbol{\Pi}_{i}^{I}[\eta]$ denotes the TPM for the $i$-th path due to vehicle movement and the CKM generates the beam TPMs $\boldsymbol{\Pi}_{i}^{II}[\eta]$ for each path based on the predicted position $\boldsymbol{\hat{q}}[\eta+1|\eta]$ in the C-Domain. Consequently, the variation in AoA associated with each path can be effectively modeled as a discrete Markov process [6].
Based on the B-Domain state evolution model and the updated B-Domain state $\boldsymbol{\hat{\theta}}[\eta]$, B-Domain state prediction provides two kinds of angle prior information, i.e., the soft prediction $\{\boldsymbol{p}(\theta_{i}[\eta+1|\eta])\}_{i=1}^{P}$ and the hard prediction $\boldsymbol{\hat{\theta}}[\eta+1|\eta]$, where the former is exploited to assist B-Domain state updates while the latter is employed for designing the predictive beamformer. Then, based on the hard prediction $\boldsymbol{\hat{\theta}}[\eta+1|\eta]$, the predictive BF matrix $\boldsymbol{F}[\eta+1|\eta]$ for the next time slot is designed to minimize the maximum AoA estimation error among multiple propagation paths, which determines the received echo signal $\boldsymbol{R}[\eta+1]$ for the next time slot. By incorporating both the soft prediction $\boldsymbol{p}(\theta_{i}[\eta+1|\eta])$ and new measurements $\boldsymbol{R}_{i}[\eta+1]$ in the $[\eta+1]$-th time slot, the B-Domain state $\boldsymbol{\hat{\theta}}[\eta+1]$ can be effectively updated.
Moreover, CKM directly links the coordinate and beam domains and embeds environmental awareness into the tracking process across both domains. From the C-Domain to the B-Domain, CKM enhances beam tracking performance via mapping the state prediction of the vehicle to a prior AoA information, particularly beneficial for NLoS paths characterized by lower sensing signal-to-noise ratios (SNRs). From the B-Domain to the C-Domain, CKM is reversely utilized to infer the state of the vehicle based on multipath parameter measurements for the LoS-absent case. As a result, continuous C-Domain tracking is maintained for both LoS-present and LoS-absent cases. Moreover, since the coherence time of the C-Domain state evolution model significantly exceeds that of the B-Domain state evolution model, C-Domain tracking further improves the continuity in B-Domain tracking via exploiting CKM, particularly scenarios involving random path birth and death, or burst change of multipath AoA.
It is noteworthy that multipath AoA parameter $\boldsymbol{\theta}$ plays different roles in the beam and coordinate domains. In the B-Domain, it serves as a state, updated by combining the echo signal $\boldsymbol{R}[\eta]$ with the soft beam state prediction $\left\{\boldsymbol{p}(\theta_{i}[\eta|\eta-1])\right\}_{i=1}^{P}$. It is expected that this prior information $\{\boldsymbol{p}(\theta_{i}[\eta|\eta-1])\}_{i=1}^{P}$ can enhance the beam state update accuracy, particularly in low SNR scenarios. In the C-Domain, $\boldsymbol{\theta}$ is treated as a measurement rather than a state, which is utilized for updating the state of the vehicle.
In one time slot, the proposed dual-domain tracking scheme alternately updates the states defined in both domains and facilitates information exchanges between them via the CKM. Across time slots, the proposed dual-domain still follows the traditional Kalman filter framework, where state prediction is performed in both domains to provide a prior information for state updates in the next time slot. Different from traditional passive tracking approaches in the field of radar, ISAC provides additional degrees of freedom for designing sensing signals for the next time slot based on the predicted AoAs of multiple paths. Consequently, a predictive BF design based on the B-Domain state prediction scheme is critical for improving the performance of the proposed dual-domain tracking.
IV CKM-assisted C-Domain Tracking
In this section, we propose a CKM-assisted coordinated domain EKF tracking scheme that is applicable for both LoS-present and LoS-absent cases.
IV-A C-Domain State Evolution Model
Considering a constant velocity movement model for the moving vehicle within the maximum time of interest [18], the C-Domain state evolution model is given by $\boldsymbol{\alpha}[\eta+1]=\boldsymbol{E}\boldsymbol{\alpha}[\eta]+\boldsymbol{w}_{\alpha}[\eta+1]$, where $\boldsymbol{E}=\left[1,0,\Delta T;0,1,0;0,0,1\right]^{T}\in\mathbb{R}^{3\times 3}$ is the state transition matrix and $\boldsymbol{w}_{\alpha}[\eta+1]=[w_{q_{x}}[\eta+1],w_{q_{y}}[\eta+1],w_{v}[\eta+1]]^{T}$ denotes the state evolution noise. Assuming mutually independent state evolution noises [3, 18], we have $\boldsymbol{w}_{\alpha}[\eta+1]\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{Q}_{\alpha})$ with $\boldsymbol{Q}_{\alpha}=\operatorname{diag}\{\sigma^{2}_{q_{x}},\sigma^{2}_{q_{y}},\sigma^{2}_{v}\}$, which can be acquired by the RSU through long-term measurements.
IV-B C-Domain Measurement Model for LoS-present Case
If the LoS path exists in the $[\eta]$-th slot, the C-Domain measurements consist of the AoA, delay, and Doppler shift of the LoS path, which is denoted with the subscript ${}^{\prime\prime}1^{\prime\prime}$. Based on the geometric relationships depicted in Fig. 1, the C-Domain measurement model is given by
$\begin{cases}&\tau_{1}[\eta]=\frac{2||\boldsymbol{q}[\eta]||_{2}}{c}+w_{\tau}[\eta]\in\mathbb{R},\\ &\mu_{1}[\eta]=-\frac{2v_{1}[\eta]q_{y}[\eta]f_{c}}{c||\boldsymbol{q}[\eta]||_{2}}+w_{\mu}[\eta]\in\mathbb{R},\\ &\cos(\theta_{1}[\eta])=\frac{q_{x}[\eta]}{||\boldsymbol{q}[\eta]||_{2}}+w_{\cos\theta}[\eta]\in(-1,1),\end{cases}$ (4)
where $c$ is the speed of light and $f_{c}$ is the signal carrier frequency. Gaussian random variables $w_{\tau}[\eta]$, $w_{\mu}[\eta]$, and $w_{\cos\theta}[\eta]$ denote the corresponding measurement noises with zero mean and variances $\sigma^{2}_{\tau}$, $\sigma^{2}_{\mu}$, and $\sigma^{2}_{\cos\theta}$, respectively. Collecting all the C-Domain observable parameters in a vector $\boldsymbol{\beta}_{1}[\eta]=[\tau_{1}[\eta],\mu_{1}[\eta],\cos(\theta_{1}[\eta])]^{T}\in\mathbb{R}^{3\times 1}$, the measurement model is given by
$\boldsymbol{\beta}_{1}[\eta]=g_{1}(\boldsymbol{\alpha}[\eta])+\boldsymbol{w}_{\beta_{1}}[\eta],$ (5)
where $\boldsymbol{w}_{\beta_{1}}[\eta]=[w_{\tau}[\eta],w_{\mu}[\eta],w_{\cos\theta}[\eta]]^{T}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{Q}_{\beta_{1}})$ and $\boldsymbol{Q}_{\beta_{1}}=\operatorname{diag}\{\sigma^{2}_{\tau},\sigma^{2}_{\mu},\sigma^{2}_{\cos\theta}\}$. The non-linear measurement function $g_{1}(\cdot):\mathbb{R}^{3\times 1}\to\mathbb{R}^{3\times 1}$ is defined by (4).
IV-C C-Domain Measurement Model for LoS-absent Case
In practice, the LoS path might be shadowed by surrounding buildings or large scatterers in neighboring lanes. In LoS-absent cases, the C-Domain measurement vector is given by $\boldsymbol{\beta}_{2}[\eta]=[\boldsymbol{\tau}[\eta];\boldsymbol{\mu}[\eta];\cos(\boldsymbol{\theta}[\eta])]\in\mathbb{C}^{3P\times 1}$, where $\boldsymbol{\tau}[\eta]=\left[\tau_{1}[\eta],\dots,\tau_{P}[\eta]\right]^{T}$, $\boldsymbol{\mu}[\eta]=\left[\mu_{1}[\eta],\dots,\mu_{P}[\eta]\right]^{T}$, and $\cos(\boldsymbol{\theta}[\eta])=\left[\cos(\theta_{1}[\eta]),\dots,\cos(\theta_{P}[\eta])\right]^{T}$. For LoS-absent cases, the measured NLoS path parameters are not directly related to the state of the vehicle. Fortunately, CKM provides the location-specific channel parameters, which maps the position and the velocity of the vehicle to the path delays, Doppler shifts, and angular information. Therefore, the CKM itself provides a measurement function which is specified as $g_{2}(\cdot):\mathbb{R}^{3\times 1}\to\mathbb{R}^{3{P}\times 1}$ that maps the state vector $\boldsymbol{\alpha}[\eta]$ to the measurement vector $\boldsymbol{\beta}_{2}[\eta]$. In other words, the measurement model is given by
$\boldsymbol{\beta}_{2}[\eta]=g_{2}(\boldsymbol{\alpha}[\eta])+\boldsymbol{w}_{\beta_{2}}[\eta],$ (6)
where $\boldsymbol{w}_{\beta_{2}}[\eta]\sim\mathcal{N}(0,\boldsymbol{Q}_{\beta_{2}})\in\mathbb{R}^{3{P}\times 1}$ is the measurement noise vector and the corresponding measurement noise covariance matrix for the LoS-absent case is assumed [^5] In this paper, we assume the same measurement noise variance for the same type of channel parameter estimation in (5) and (6) as the channel parameters stored in the CKM come from long-term channel measurement as well [12]. as $\boldsymbol{Q}_{\beta_{2}}=\operatorname{diag}\{\sigma_{\tau}^{2}\boldsymbol{I}_{P},\sigma_{\mu}^{2}\boldsymbol{I}_{P},\sigma_{\cos\theta}^{2}\boldsymbol{I}_{P}\}$. As commonly adopted in literature [18, 9], we assume the estimation errors of C-Domain measurement to be Gaussian distributed, as required by the EKF framework. In practice, $\sigma_{\tau}^{2}$, $\sigma_{\cos\theta}^{2}$, and $\sigma_{\cos\theta}^{2}$ depend on antenna aperture, bandwidth, frame length, and even BF strategy. The relationship between path parameter estimation errors and those resources above is specifically determined by parameter estimation methods. Therefore, for simplicity, this paper assumes fixed measurement noise variances in the adopted C-Domain.
IV-D EKF-based C-Domain Tracking
Due to the non-linear measurement model for both the LoS-present case in (5) and LoS-absent case in (6), we adopt the EKF [34] to track the state of the vehicle in the C-Domain. In the $\eta$-slot, assuming we have the knowledge of the state vector ${\boldsymbol{\alpha}}[\eta]\in\mathbb{R}^{3\times 1}$ estimated at the $\eta$-slot, the state prediction for the $[\eta+1]$-slot is given by
$\hat{\boldsymbol{\alpha}}[\eta+1|\eta]=\boldsymbol{E}{\boldsymbol{\alpha}}[\eta],$ (7)
and the prediction covariance matrix $\boldsymbol{C}_{r}[\eta+1|\eta]\in\mathbb{R}^{3\times 3}$, $r=1,2$ is given by $\boldsymbol{C}_{r}[\eta+1|\eta]=\boldsymbol{E}\boldsymbol{C}_{r}[\eta]\boldsymbol{E}^{T}+\boldsymbol{Q}_{\alpha}$, where $\boldsymbol{C}_{r}[\eta]$ is the posterior covariance matrix for the C-Domain state update in the $[\eta]$-th slot.
IV-D1 LoS-present case
To update the state in the $[\eta]$-th time slot, the EKF employs a linearized measurement model around the predicted state [^6] In our notation, only the state prediction and state update are denoted with a hat (e.g., $\hat{\boldsymbol{\alpha}}[\eta|\eta-1]$ and $\hat{\boldsymbol{\alpha}}[\eta]$); other variables, such as the measurement $\boldsymbol{\beta}[\eta|\eta-1]$ and $\boldsymbol{\beta}[\eta]$, are presented without a hat.. Specifically, we have: $\boldsymbol{\beta}_{1}[\eta]\approx g_{1}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{G}_{1}(\boldsymbol{\alpha}[\eta]-\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{w}_{\beta_{1}}[\eta]$, where $\hat{\boldsymbol{\alpha}}[\eta|\eta-1]$ is the predicted C-Domain state in $[\eta-1]$-th time slot, $\boldsymbol{G}_{1}\in\mathbb{R}^{3\times 3}$ represents the Jacobian matrix of $g_{1}$ with respect to the state vector and it is given by (8)
$\boldsymbol{G}_{1}$ $=\left.\frac{\partial{g}_{1}}{\partial\boldsymbol{\alpha}}\right|_{\hat{\boldsymbol{\alpha}}[\eta|\eta-1]}=\left.\begin{bmatrix}\frac{\partial\tau_{1}[\eta]}{\partial q_{x}[\eta]}&\frac{\partial\tau_{1}[\eta]}{\partial q_{y}[\eta]}&0\\ \frac{\partial\mu_{1}[\eta]}{\partial q_{x}[\eta]}&\frac{\partial\mu_{1}[\eta]}{\partial q_{y}[\eta]}&\frac{\partial\mu_{1}[\eta]}{\partial v_{1}[\eta]}\\ \frac{\partial\cos(\theta_{1}[\eta])}{\partial q_{x}[\eta]}&\frac{\partial\cos(\theta_{1}[\eta])}{\partial q_{y}[\eta]}&0\end{bmatrix}\right|_{\hat{\boldsymbol{\alpha}}[\eta|\eta-1]}$ (8)
$=\left.\begin{bmatrix}\frac{2q_{x}[\eta]}{c||\boldsymbol{q}[\eta]||_{2}}&\frac{2q_{y}[\eta]}{c||\boldsymbol{q}[\eta]||_{2}}&0\\ -\frac{2v_{1}[\eta]f_{c}q_{y}[\eta]q_{x}[\eta]}{c||\boldsymbol{q}[\eta]||_{2}^{3}}&\frac{2f_{c}v_{1}[\eta](q_{x}[\eta])^{2}}{c||\boldsymbol{q}[\eta]||_{2}^{3}}&-\frac{2q_{y}[\eta]f_{c}}{c||\boldsymbol{q}[\eta]||_{2}}\\ \frac{q_{y}[\eta]^{2}}{||\boldsymbol{q}[\eta]||_{2}^{3}}&-\frac{q_{x}[\eta]q_{y}[\eta]}{||\boldsymbol{q}[\eta]||_{2}^{3}}&0\end{bmatrix}\right|_{\hat{\boldsymbol{\alpha}}[\eta|\eta-1]}.$
By substituting the state prediction $\hat{\boldsymbol{\alpha}}[\eta|\eta-1]$ into the equation above, $\boldsymbol{G}_{1}$ is obtained.
IV-D2 LoS-absent case
In the LoS-absent case, the measurement function $g_{2}(\cdot)$ in (6) leveraging CKM provides a look-up table for obtaining multipath parameter measurements based on the location of the vehicle. To update the state of the vehicle, a linearized measurement model can be established:
${\boldsymbol{\beta}_{2}}[\eta]\approx g_{2}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{G}_{2}(\boldsymbol{\alpha}[\eta]-\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{w}_{\beta_{2}}[\eta],$ (9)
where $g_{2}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])\in\mathbb{R}^{3P\times 1}$ can be obtained via referring to CKM based on the predicted vehicle state. The Jacobian matrix $\boldsymbol{G}_{2}$ represents the partial derivatives of the measurement function $g_{2}(\cdot)$ with respect to the C-Domain state vector and is computed numerically as follow:
$\boldsymbol{G}_{2}$ $=\left.\frac{\partial g_{2}}{\partial\boldsymbol{\alpha}}\right|_{\hat{\boldsymbol{\alpha}}[\eta|\eta-1]}=\left.[\frac{\partial g_{2}}{\partial q_{x}},\frac{\partial g_{2}}{\partial q_{y}},\frac{\partial g_{2}}{\partial v_{1}}]\right|_{\hat{\boldsymbol{\alpha}}[\eta|\eta-1]},$ (10)
where $\frac{\partial g_{2}}{\partial q_{x}}={[g_{2}(q_{x}[\eta]+\Delta q_{x},q_{y}[\eta],v_{1}[\eta])-g_{2}(\boldsymbol{\alpha}[\eta])]}/{\Delta q_{x}}$, $\frac{\partial g_{2}}{\partial q_{y}}={[(g_{2}(q_{x}[\eta],q_{y}[\eta]+\Delta q_{y},v_{1}[\eta])-g_{2}(\boldsymbol{\alpha}[\eta])]}/{\Delta q_{y}}$, and $\frac{\partial g_{2}}{\partial v_{1}}={[(g_{2}(q_{x}[\eta],q_{y}[\eta],v_{1}[\eta]+\Delta v_{1})-g_{2}(\boldsymbol{\alpha}[\eta])]}/{\Delta v_{1}}$, $\Delta q_{x}$, $\Delta q_{y}$, and $\Delta v_{1}$ denote small perturbations in the C-Domain state. Note that each element of $\boldsymbol{G}_{2}$ quantifies how the multipath parameters change with respect to the C-Domain state vector.
Following the standard EKF framework [18], the procedures for proposed C-Domain tracking in both LoS-present $(r=1)$ and LoS-absent $(r=2)$ scenarios are outlined below: (1) Linearization: $\boldsymbol{\beta}_{r}[\eta]\approx g_{r}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{G}_{r}(\boldsymbol{\alpha}[\eta]-\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{w}_{\beta_{r}}[\eta]$, (2) Kalman gain matrix: $\boldsymbol{K}_{r}=\boldsymbol{C}_{r}[\eta|\eta-1]\boldsymbol{G}_{r}^{T}(\boldsymbol{G}_{r}\boldsymbol{C}_{r}[\eta|\eta-1]\boldsymbol{G}_{r}^{T}+\boldsymbol{Q}_{\beta_{r}})^{-1}$, (3) State update: $\hat{\boldsymbol{\alpha}}[\eta]=\hat{\boldsymbol{\alpha}}[\eta|\eta-1]+\boldsymbol{K}_{r}(\boldsymbol{\beta}_{r}[\eta]-g_{r}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1]))$, (4) Posterior covariance matrix: $\boldsymbol{C}_{r}[\eta]=(\boldsymbol{I}_{3}-\boldsymbol{K}_{r}\boldsymbol{G}_{r})\boldsymbol{C}_{r}[\eta|\eta-1]$, (5) State prediction: $\hat{\boldsymbol{\alpha}}[\eta+1|\eta]=\boldsymbol{E}\hat{\boldsymbol{\alpha}}[\eta]$, (6) Prediction covariance matrix: $\boldsymbol{C}_{r}[\eta+1|\eta]=\boldsymbol{E}\boldsymbol{C}_{r}[\eta]\boldsymbol{E}^{T}+\boldsymbol{Q}_{\alpha}$. Interested readers can refer to [18, 9, 31, 3] for further details of EKF. The proposed EKF scheme requires identifying whether the LoS is present or not, a process already proposed in [4] and [9]. The computational complexity of the proposed EKF scheme is constant, i.e., $\mathcal{O}(1)$, in the LoS-present case, whereas in the LoS-absent case, it increases to $\mathcal{O}(P^{3})$ due to the inversion of a $3P\times 3P$ matrix required in the Kalman gain computation.
V CKM-assisted B-Domain Tracking
V-A B-Domain State Evolution Model
Similar to [35] and [6], define $\Theta\triangleq\{\tilde{\theta}_{\kappa}={\pi(\kappa-1)}/{N_{\theta}}\}^{N_{\theta}}_{\kappa=1}$ as an angular grid with ${N}_{\theta}$ equal segments spanning the entire angular space $[0,\pi)$. Also, the B-Domain state evolution model is defined as a discrete Markov process, i.e.,
$\small\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])=\boldsymbol{p}(\theta_{i}[\eta])\boldsymbol{\Pi}_{i}[\eta]\in[0,1]^{1\times N_{\theta}},$ (11)
where $\boldsymbol{\Pi}_{i}[\eta]\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ denotes the TPM for the AoA of $i$-th path in the $[\eta]$-th time slot, $\boldsymbol{p}({\theta}_{i}[\eta])$ and $\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])$ denote the probability mass function (PMF) of the AoA for the $i$-th path in the $[\eta]$-th slot and the predicted PMF of the AoA of $i$-th path in the $[\eta+1]$-th time slot, respectively. The $(\kappa,\iota)$-th element of the TPM for the $i$-th path in the $[\eta]$-th slot is given by $\boldsymbol{\Pi}_{i}[\eta]_{\kappa,\iota}=\operatorname{Pr}\{\theta_{i}[\eta+1]=\tilde{\theta}_{\iota}\mid\theta_{i}[\eta]=\tilde{\theta}_{\kappa}\},\forall\kappa,\iota\in\{1,\dots,N_{\theta}\}$, where $\operatorname{Pr}\{\theta_{i}[\eta+1]=\tilde{\theta}_{\iota}\mid\theta_{i}[\eta]=\tilde{\theta}_{\kappa}\}$ denotes the transition probability of the AoA of the $i$-th path from angle $\tilde{\theta}_{\kappa}$ at time slot $[\eta]$ to $\tilde{\theta}_{\iota}$ at time slot $[\eta+1]$. Clearly, the TPM should satisfy the following requirements: (1) ${0}\leq\boldsymbol{\Pi}_{i}[\eta]_{\kappa,\iota}\leq{1}$ for $\forall\kappa,\iota\in\{1,\dots,N_{\theta}\}$. (2) The sum of each row of $\boldsymbol{\Pi}_{i}[\eta]$ should be equal to 1, i.e., $||\boldsymbol{\Pi}_{i}[\eta]_{\kappa,:}||_{1}=1,\forall\kappa\in\{1,\dots,N_{\theta}\}$.
Both the channel temporal correlation and the CKM provide a prior information on beam transition and thus affect the TPM. Therefore, the angular TPM for the $i$-th path at $[\eta]$-th slot is modeled as:
$\small\boldsymbol{\Pi}_{i}[\eta]=(1-c_{\Pi}[\eta])\cdot\boldsymbol{\Pi}_{i}^{I}[\eta]+c_{\Pi}[\eta]\cdot\boldsymbol{\Pi}_{i}^{II}[\eta],$ (12)
where $\boldsymbol{\Pi}_{i}^{I}[\eta]$ denotes the TPM caused by the channel temporal correlation and $\boldsymbol{\Pi}_{i}^{II}[\eta]$ denotes the TPM designed using the channel a priori information provided by CKM. Real constant $c_{\Pi}[\eta]\in[0,1]$ is a weight parameter that is tuned to characterize the contribution of a prior information provided by channel temporal correlation and CKM. The value of $c_{\Pi}[\eta]$ depends on the condition of beam transition, the received echo SNR, and the horizontal speed of the target vehicle relative to the RSU.
As mentioned earlier, the AoA of the $i$-th path between two consecutive time slots should remain close even in high-mobility wireless networks, provided there is no random birth/death of paths. Following [35] and [36], the $(\kappa,\iota)$-th element of $\boldsymbol{\Pi}_{i}^{I}$ of a path in the $[\eta]$-th slot is given by
$\small\left\{\boldsymbol{\Pi}_{i}^{I}[\eta]\right\}_{\kappa,\iota}=\begin{cases}\zeta\xi^{\left|\kappa-\iota\right|},&{\text{if}}\left|\kappa-\iota\right|\leq\epsilon,\\ {0},&{\text{otherwise,}}\end{cases}$ (13)
where $\forall i\in\{1,\dots,P\}$. The parameter $\xi\in[0,1]$ characterizes the variation speed of angles over time, where $\xi\rightarrow 0$ indicates almost no change and $\xi\rightarrow 1$ indicates significant change. The variable $\epsilon$ represents the boundary of AoA variation and $\zeta$ is a normalization coefficient as in [17, 6]. Both $\xi$ and $\epsilon$ depend on the relative speed of the vehicle to the RSU, given by $\sin(\hat{\theta}_{1}[\eta+1|\eta])\hat{v}_{1}[\eta+1|\eta]$. On the other hand, the $(\kappa,\iota)$-th element of $\boldsymbol{\Pi}_{i}^{II}$ in the $[\eta]$-th slot is given by
$\small\left\{\boldsymbol{\Pi}_{i}^{II}[\eta]\right\}_{\kappa,\iota}=\varsigma\exp\left(-\frac{(\hat{\theta}_{i}[\eta+1|\eta]-\tilde{\theta_{\iota}})^{2}}{2\sigma_{\text{CKM}}^{2}}\right),\forall\kappa,$ (14)
where $\varsigma=1/\left(\sum_{\iota=1}^{N_{\theta}}\exp\left(-\frac{(\hat{\theta}_{i}[\eta+1|\eta]-\tilde{\theta_{\iota}})^{2}}{2\sigma_{\text{CKM}}^{2}}\right)\right)$ is the normalization coefficient, $\hat{\theta}_{i}[\eta+1|\eta]$ is the predicted AoA of the $i$-th path provided by CKM in (3) based on the predicted vehicle state $\hat{\boldsymbol{\alpha}}[\eta+1|\eta]$ in (7). $\tilde{\theta}_{\iota}=\pi(\iota-1)/N_{\theta}$ denotes the $\iota$-th angle on the grid $\Theta$ and $\sigma_{\text{CKM}}^{2}$ serves as a variance of beam transition accounting for noise in the angle measurements and uncertainty in the CKM. A typical example of $\boldsymbol{\Pi}_{i}$, $\boldsymbol{\Pi}_{i}^{I}$, and $\boldsymbol{\Pi}_{i}^{II}$ are illustrated in Fig. 3.
Figure 3: Angular TPM for the $i$-th path.
Moreover, the B-Domain state evolution model should depend on the condition of beam transition.
Figure 4: Angular variation over time for each path.
In Fig.4, the angular variations of two paths over time are demonstrated, where path 1 represents a LoS path and path 2 represents an NLoS path. During the time of interest, the condition of beam transition can be established according to the following three types of angular variations:
  • Stationary Transition $(I)$: In this case, both $\boldsymbol{\Pi}_{i}^{I}$ and $\boldsymbol{\Pi}_{i}^{II}$ are available due to the time-correlated variations in AoA and the CKM mapping between the position of vehicle and AoA. In this case, $c_{\Pi}\in(0,1)$.
  • Nonstationary but Predictable Transition $(II)$: The angle may abruptly change due to the fixed scatterers in the environment which have been recorded in the CKM (shown by the yellow arrows in the figure). In this case, the angle transition is not time-correlated and $\boldsymbol{\Pi}_{i}^{I}$ can not be exploited. Fortunately, CKM has stored this beam transition and thus the angle transition is still predictable based on $\boldsymbol{\Pi}_{i}^{II}$. In this case $c_{\Pi}$ is equal to 1.
  • Nonstationary and Unpredictable Transition $(III)$: The random block is caused by randomly appearing scatterers/obstacles in the environment (illustrated by the red asterisks in the figure). For this case, AoA prediction is impossible and no a priori information can be used. In this case, a confirm distribution can be adopted, i.e., $\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])=\frac{1}{N_{\theta}}\cdot\boldsymbol{1}_{N_{\theta}}$.
The hypothesis testing for LoS path blockage has been proposed in [4] and [9] for improving target tracking performance. Similarly, based on the estimated delay and Doppler shift between consecutive time slots, the RSU can determine the condition of beam transition for each path. In this paper, we assume that the hypothesis testing for cases of the angular variation model is perfect for simplicity, which facilitates the B-Domain state evolution model establishment in the proposed dual-domain tracking and predictive beam forming framework.
V-B B-Domain Measurement Model
After receiving the reflected echo, the RSU can calculate the cross-ambiguity function (CAF) [32] between the received signal $\boldsymbol{R}$ and the transmitted signal $\tilde{\boldsymbol{S}}$ to estimate the delay and Doppler shift of multiple paths. Following (2), a peak occurs at the matched filter output when $\tau$ and $\mu$ are perfectly tuned to the corresponding delay and Doppler shifts of the $i$-th path, respectively, i.e., $\{\hat{l}_{i},\hat{k}_{i}\}={\text{arg max}}_{l_{i},k_{i}}||\boldsymbol{R}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}||_{F}^{2},i=1,\dots,P$.
Lemma 1
Assuming $L\to\infty$ and $\frac{1}{T_{p}}\to\infty$, having obtained the perfect delay and Doppler shift estimates, the received signal $\boldsymbol{R}_{i}$ associated with the $i$-th path can be separated and is given by
$\boldsymbol{R}_{i}\approx L\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}+\boldsymbol{Z}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}\in\mathbb{C}^{N_{r}\times N_{s}},$ (15)
for any $i=1,\cdots,P$.
Proof:
Substitute (2) into the received signal after matched filtering:
$\small\boldsymbol{R}_{i}=\underbrace{\sum_{j=1}^{P}\beta_{j}\boldsymbol{b}(\theta_{j})\boldsymbol{a}^{H}(\theta_{j})\boldsymbol{F}{\boldsymbol{S}_{l_{j}}}{\boldsymbol{\Lambda}_{k_{j}}}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}}_{\text{Signal and interference: }\boldsymbol{R}_{i}^{\prime}}+\underbrace{\boldsymbol{Z}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}}_{\text{Noise: }\boldsymbol{R}_{i}^{\prime\prime}},$ (16)
where $\boldsymbol{R}_{i}^{\prime}$ can be approximated as $\boldsymbol{R}_{i}^{\prime}\approx L\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}$ based on the assumptions above. ∎
Lemma 1 enables the RSU to separate the received signal from different paths for AoA estimation using the estimated delay and Doppler shift. According to (15), the received signal associated with the $i$-th path is rewritten as: $\boldsymbol{r}_{i}\triangleq{\text{vec}}(\boldsymbol{R}_{i})=L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i})+(\boldsymbol{S}_{l_{i}}^{*}\otimes\boldsymbol{Z}){\text{vec}}(\boldsymbol{\Lambda}_{k_{i}}^{H})\in\mathbb{C}^{N_{s}N_{r}\times 1}$. Given the AoA of the $i$-th path $\theta_{i}$, we can prove that $\boldsymbol{r}_{i}\sim\mathcal{CN}(L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i}),\sigma_{z}^{2}L\boldsymbol{I}_{N_{s}N_{r}})$ and the likelihood function of $\theta_{i}$ is given by
$\small p\left(\boldsymbol{r}_{i}|\theta_{i}\right)=\frac{1}{(L\pi\sigma_{z}^{2})^{N_{s}N_{r}}}\exp\{-\frac{1}{2\sigma_{z}^{2}}\bar{\boldsymbol{r}}_{i}^{H}\bar{\boldsymbol{r}}_{i}\}$ (17)
where $\bar{\boldsymbol{r}}_{i}=\boldsymbol{r}_{i}-L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i})$.
V-C B-Domain Tracking
B-Domain tracking also comprises two components, i.e., B-Domain state prediction and B-Domain state update, where the state prediction is categorized into hard and soft predictions. In time slot $\eta$, the B-Domain soft state prediction has been defined in (11). Then, the hard state prediction is given by
$\small\hat{\theta}_{i}[\eta+1|\eta]=\arg\max\operatorname{Pr}(\theta_{i}[\eta+1|\eta]).$ (18)
Based on the B-Domain soft state prediction $\boldsymbol{p}({\theta}_{i}[\eta|\eta-1])$ in the $[\eta-1]$-th time slot and new measurement $\boldsymbol{R}_{i}[\eta]$ in (15), we can update the AoA of the $i$-th path in $[\eta]$-th time slot employing the maximum a posteriori (MAP) principle, i.e.,
$\small\hat{\theta}_{i}[\eta]=\max_{\theta_{i}[\eta]}p(\theta_{i}[\eta]|\boldsymbol{R}_{i}[\eta])\propto\max_{\theta_{i}[\eta]}p\left(\boldsymbol{r}_{i}[\eta]|\theta_{i}[\eta]\right)p(\theta_{i}[\eta|\eta-1]),$ (19)
where $p(\theta_{i}[\eta|\eta-1])=\frac{1}{N_{\theta}}$ for the non-stationary and unpredicatable transition $(III)$ case and (19) degenerates to the maximum likelihood estimation. Note that AoA estimation has been extensively studied in the field of both signal processing and wireless communication and various well-developed algorithms can be employed to solve (19), such as the subspace approach [37] and compressed sensing techniques [38]. In this work, we focus more on the dual-domain tracking framework and thus simply adopt an exhaustive search method to solve (19) as only one angular parameter is involved in (19).
VI CKM-assisted Predictive BF
In this section, we first derive the CRB of multipath AoA estimation and then design predictive BF for the subsequent time slot to minimize the maximum CRB for AoA estimation of each path, exploiting the channel a priori information obtained from the previous time slot. It is worth noting that accurate beam tracking benefits both target tracking and communication, which motivates us to design predictive BF for optimizing the beam tracking performance in the next time slot. The trade-off between dual-domain tracking and communication performance is a critical issue of the considered system. However, C-Domain tracking performance depends on B-Domain tracking performance and thus could be a complicated function of predictive BF strategy. Moreover, optimizing communication performance needs to consider the uncertainty of predicted CSI based on the predicted AoA for each path, which would lead to a complicated optimization problem. As a first attempt to propose a dual-domain tracking framework, this paper only considers the beam tracking performance in predictive BF design and the optimization of the trade-off between sensing and communication in the context of dual-domain tracking is left for our future work.
VI-A CRB for Multi-path AoA Estimation
CRB serves as a lower bound for the variances of any unbiased estimators [34], which can characterize the multipath AoAs estimation performance. For each path, given the likelihood function in (17), the Fisher information (FI) [34] of unknown parameter $\theta_{i}$ is given by $\mathcal{J}\left(\theta_{i}\right)=-\mathbb{E}\left(\frac{\partial^{2}\ln p\left(\boldsymbol{r}_{i}\mid\theta_{i}\right)}{\partial\theta_{i}^{2}}\right)=\frac{L^{2}|\beta_{i}|^{2}}{\sigma_{z}^{2}}\boldsymbol{j}(\theta_{i})^{H}\boldsymbol{j}(\theta_{i})$, where $\boldsymbol{j}(\theta_{i})=(\boldsymbol{F}^{T}\otimes\frac{\partial\boldsymbol{b}(\theta_{i})}{\partial\theta_{i}})\boldsymbol{a}^{*}(\theta_{i})$, and the derivative of the steering vector $\boldsymbol{b}(\theta_{i})$ with respect to $\theta_{i}$ is given by ${\partial\boldsymbol{b}(\theta_{i})}/{\partial\theta_{i}}=j\pi\cos(\theta_{i}-{\pi}/{2})\operatorname{diag}(0,\ldots,N_{r}-1)\boldsymbol{b}(\theta_{i})$. Based on the definition of CRB [34], the mean squared error (MSE) of AoA estimation is lower bounded by the inverse of $\mathcal{J}\left(\theta_{i}\right)$, which is given by
$\small\mathbb{E}[(\hat{\theta}_{i}-\theta_{i})^{2}]\geq{1}/{\mathcal{J}(\theta_{i})}\triangleq\operatorname{CRB}\left(\theta_{i}\right).$ (20)
Remark 1: The posterior Cramér-Rao bound (PCRB), which incorporates both the measured data and the a priori information of $\theta_{i}$, is commonly used to characterize the estimation performance in beam tracking frameworks [18]. However, as discussed in Fig. 4, a priori information might be unavailable for the third case of the condition of beam transition, i.e., non-stationary and unpredictable transition. Moreover, even for the first two types of AoA variation, the prior information obtained from TPM is numerical and thus cannot be integrated into PCRB derivation. Consequently, we adopt CRB as multipath AoA estimation performance metric. Note that in the $[\eta]$-th time slot, CRB in (20) depends on ground truth value of $\theta_{i}$ in the $[\eta+1]$-th time slot, which is unavailable. However, the B-Domain hard state prediction in (18) provides an AoA prediction $\hat{\theta}_{i}[\eta+1|\eta]$ which can be substituted into (20) to compute CRB. We note that this is one way to embed a prior information of $\theta_{i}$ into its predictive estimation error, as commonly adopted in the literature [18].
VI-B Predictive BF Design
In $[\eta]$-th time slot, the predictive BF design to minimize the maximum predictive AoA estimation error $\mathrm{CRB}(\theta_{i}[\eta+1|\eta])$ is formulated as the following optimization problem
$\small\mathop{\mathrm{minimize}}_{\boldsymbol{F}[\eta+1|\eta]}\max_{i}\ \{\mathrm{CRB}(\theta_{i}[\eta+1|\eta])\}\ \text{s.t.}||\boldsymbol{F}[\eta+1|\eta]||_{F}^{2}\leq P_{\text{t}}.$ (21)
where $\boldsymbol{F}[\eta+1|\eta]\in\mathbb{C}^{N_{t}\times N_{s}}$ is the predictive BF matrix and $P_{\text{t}}$ is the total transmit power of the RSU. Utilizing (20), the predictive BF design problem can be rewritten as:
$\small\mathop{\mathrm{maximize}}_{\boldsymbol{F}[\eta+1|\eta]}\min_{i}\ \{\mathcal{J}(\theta_{i}[\eta+1|\eta])\}\ \text{s.t. }\left\|\boldsymbol{F}[\eta+1|\eta]\right\|_{F}^{2}\leq P_{\text{t}}.$ (22)
Note that problem (22) is difficult to be solved globally due to the non-convex objective function. As a compromise, we propose a two-step optimization scheme with beam selection and multi-beam power allocation. In particular, the predictive BF matrix $\boldsymbol{F}[\eta+1|\eta]$ is defined as the product of an analog beam steering matrix $\boldsymbol{A}[\eta+1|\eta]\in\mathbb{C}^{N_{t}\times N_{s}}$ and a diagonal power allocation matrix $\boldsymbol{\Gamma}[\eta+1|\eta]\in\mathbb{R}^{N_{s}\times N_{s}}$, i.e., $\boldsymbol{F}[\eta+1|\eta]=\boldsymbol{A}[\eta+1|\eta]\boldsymbol{\Gamma}[\eta+1|\eta]$, where $\boldsymbol{\Gamma}[\eta+1|\eta]\triangleq\operatorname{diag}(\sqrt{\gamma_{1}[\eta+1|\eta]},\cdots,\sqrt{\gamma_{N_{s}}[\eta+1|\eta]})$, and $\gamma_{n_{s}}[\eta+1|\eta]$ denotes the power allocated to the $n_{s}$-th beam in the $[\eta+1]$-th slot.
VI-B1 Beam Selection
The steering matrix $\boldsymbol{A}[\eta+1|\eta]$ is designed based on the B-Domain hard state prediction and its $n_{s}$-th column is given by
$\boldsymbol{A}_{:,n_{s}}[\eta+1|\eta]=\boldsymbol{a}(\hat{\theta}_{n_{s}}[\eta+1|\eta])\in\mathbb{C}^{N_{t}\times 1},\forall n_{s},$ (23)
where $\hat{\theta}_{n_{s}}[\eta+1|\eta]$ is the predicted AoA for the $n_{s}$-th path at the $[\eta]$-th time slot in (18) [^7] If $N_{s}<P$, the predicted AoAs with higher channel path gains should be selected for BF. Conversely, if $N_{s}\geq P$, the predicted AoAs for all $P$ paths can be selected. .
VI-B2 Power Allocation
The second step is to acquire the optimal power allocation matrix $\boldsymbol{\Gamma}[\eta+1|\eta]$ based on the given steering matrix $\boldsymbol{A}[\eta+1|\eta]$. The objective function of (22) given $\boldsymbol{A}[\eta+1|\eta]$ is given by (24) at the top of next page, where $a_{i}\triangleq\beta_{i}\sin\left(\theta_{i}\right),\forall i$ for simplicity.
$\mathcal{J}(\theta_{i}[\eta+1|\eta])$ $=\frac{L^{2}\pi^{2}|a_{i}[\eta+1|\eta]|^{2}}{\sigma_{z}^{2}}\left\{\left[\left(\boldsymbol{\Gamma}^{T}[\eta+1|\eta]\boldsymbol{A}^{T}[\eta+1|\eta]\right)\otimes\left(\operatorname{diag}(0,\ldots,N_{r}-1)\boldsymbol{b}(\theta_{i}[\eta+1|\eta])\right)\right]\boldsymbol{a}^{*}(\theta_{i}[\eta+1|\eta])\right\}^{H}$ (24)
$\quad\cdot\left\{\left[\left(\boldsymbol{\Gamma}^{T}[\eta+1|\eta]\boldsymbol{A}^{T}[\eta+1|\eta]\right)\otimes\left(\operatorname{diag}(0,\ldots,N_{r}-1)\boldsymbol{b}(\theta_{i}[\eta+1|\eta])\right)\right]\boldsymbol{a}^{*}(\theta_{i}[\eta+1|\eta])\right\}.$
Since we focus on the power allocation for each time slot, the time index in (24) is omitted, i.e., $\beta_{i}[\eta+1|\eta]$, $\boldsymbol{\Gamma}[\eta+1|\eta]$, $\boldsymbol{A}[\eta+1|\eta]$, $\theta_{i}[\eta+1|\eta]$ and $a_{i}[\eta+1|\eta]$ are represented by $\beta_{i}$, $\boldsymbol{\Gamma}$, $\boldsymbol{A}$, $\theta_{i}$ and $a_{i}$, respectively. Using the mixed-product property of the Kronecker product, we can obtain $\mathcal{J}(\theta_{i})=({L^{2}\pi^{2}a_{i}^{2}}/{\sigma_{z}^{2}})\boldsymbol{c}^{H}(\theta_{i})(\boldsymbol{\Gamma}^{T}\otimes\operatorname{diag}(0,\ldots,N_{r}-1))^{H}(\boldsymbol{\Gamma}^{T}\otimes\operatorname{diag}(0,\ldots,N_{r}-1))\boldsymbol{c}(\theta_{i})$, where $\boldsymbol{c}(\theta_{i})\triangleq(\boldsymbol{A}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i})\in\mathbb{C}^{N_{s}N_{r}\times 1}$ can be calculated based on (23). Applying the mixed-product property of the Kronecker product again, yields $\mathcal{J}(\theta_{i})=({L^{2}\pi^{2}a_{i}^{2}}/{\sigma_{z}^{2}})\boldsymbol{c}^{H}(\theta_{i})[\boldsymbol{\Gamma}\boldsymbol{\Gamma}^{H}\otimes\operatorname{diag}^{2}(0,\ldots,N_{r}-1)]\boldsymbol{c}(\theta_{i})$. Ultimately, $\mathcal{J}\left(\theta_{i}\right)$ in (24) can be rewritten as $\mathcal{J}\left(\theta_{i}\right)=({L^{2}\pi^{2}a_{i}^{2}}/{\sigma_{z}^{2}})\sum_{{n_{s}}=1}^{N_{s}}\gamma_{n_{s}}\sum_{g=1}^{N_{r}}|\{\boldsymbol{c}(\theta_{i})\}_{({n_{s}}-1)N_{r}+g}|^{2}(g-1)^{2}$, where $\{\boldsymbol{c}(\theta_{i})\}_{k}$ denotes the $k$-th element of the vector $\boldsymbol{c}(\theta_{i})$. Further introducing auxiliary variables $t\geq 0$ to bound the minimum value of $\mathcal{J}(\theta_{i})$, problem (22) can be equivalently rewritten as
$\min_{t,\gamma_{1},\ldots,\gamma_{N_{s}}}$ $t$ (25)
s.t. $t\leq a_{i}^{2}\sum_{{n_{s}}=1}^{N_{s}}\gamma_{n_{s}}c_{i,{n_{s}}},\sum_{{n_{s}}=1}^{N_{s}}\gamma_{n_{s}}\leq\frac{P_{\text{t}}}{N_{t}},\gamma_{n_{s}}\geq 0,\forall i,\forall{n_{s}},$
where $c_{i,{n_{s}}}=\sum_{g=1}^{N_{r}}|\{\boldsymbol{c}(\theta_{i})\}_{({n_{s}}-1)N_{r}+g}|^{2}(g-1)^{2}$. After the proposed series of transformations, it can be verified that the power allocation problem (25) is convex and can be solved using off-the-shelf numerical convex programming tools, such as CVX [39].
VII Simulation Results
In this section, we present the numerical results to evaluate the performance of the proposed CKM-assisted dual-domain tracking and predictive BF framework.
VII-A Simulation Parameters And Baseline Schemes
The RSU is positioned at the origin and the vehicle starts from the coordinates $[-20\mathrm{\leavevmode\nobreak\ m},10\mathrm{\leavevmode\nobreak\ m}]^{T}$ and moves along the road with an initial speed of $10\mathrm{\leavevmode\nobreak\ m/s}$. The time of interest is $T_{\max}=4\mathrm{\leavevmode\nobreak\ s}$ with a slot duration of $\Delta T=0.02\mathrm{\leavevmode\nobreak\ s}$. The carrier frequency $f_{c}$ is set to $30\mathrm{\leavevmode\nobreak\ GHz}$ and the sample interval $T_{p}$ is $10^{-8}$ seconds. Both the maximum delay and Doppler shift indices are set as $100$. The total power transmit $P_{\text{t}}$ is $16\mathrm{\leavevmode\nobreak\ W}$ and the angular grid size is set as $N_{\theta}=7200$. For the C-Domain state evolution noises, we set the standard deviations for position $\sigma_{q_{x}},\sigma_{q_{y}}$ are $10^{-3}\mathrm{\leavevmode\nobreak\ m}$, while $\sigma_{v}$ is $10^{-3}\mathrm{\leavevmode\nobreak\ m/s}$. The standard deviations for the delay and Doppler shift measurement noise are $\sigma_{\tau}=10^{-8}\mathrm{\leavevmode\nobreak\ s}$ and $\sigma_{\mu}=20\mathrm{\leavevmode\nobreak\ Hz}$, respectively. The standard deviation of angle measurement noise is $\sigma_{\cos(\theta)}=0.01$. For the stationary transition case, the TPM constant is $c_{\Pi}=0.3$ for simplicity, and AoA variation model parameters are $\xi=0.8$, $\epsilon=\hat{v}\sin(\hat{\theta}_{1})N_{\theta}/20$, and $\sigma_{\text{CKM}}=10^{-3}$.
The ray tracing toolbox in MATLAB is leveraged to generate the ground-truth channel information [40]. A CKM is designed to record the $P=2$ strongest channel paths, denoted as path 1 and path 2, respectively. The CKM is constructed using the inverse distance weighting (IDW) of the $K$ nearest neighbors (KNN) approach with a limited set of 3200 ray tracing data samples [41]. The static shadowing caused by static scatterers in the propagation environment is assumed to emerge between the $140$-th and $175$-th time slots, whereas the random shadowing caused by a random obstacle appears with a certain blockage probability. Note that NLoS path is more likely to be blocked than LoS path in practice [42]. Therefore, we can define the blockage probability of the LoS path as $p_{\text{blk}}$ and the block probability of an NloS path as $1-(1-p_{\text{blk}})^{2}$ since an NLoS path can be treated as the concatenation of two LoS paths.
In this section, a baseline method described in [4] is considered to demonstrate the benefits of the proposed framework. In particular, the baseline algorithm in [4] integrates the result of LoS path blockage detection with an EKF-based tracking framework where the position, speed, and LoS path AoA are jointly tracked. Note that the baseline is designed for solely based on the LoS channel model and fails to account for the information of NLoS paths. Therefore, we assume that the baseline scheme can perfectly separate the received signal of LoS path in the LoS-present case, which is utilized for tracking. If the LoS path is absent, the sensing receiver ignores the observation of the NLoS path and performs state prediction only to predict the position, speed, and LoS path AoA of the vehicle as considered in [4].
VII-B Tracking Performance
Figure 5: Tracking performance for the proposed dual-domain framework versus the time index with $\sigma_{z}^{2}=10^{-9}\mathrm{\leavevmode\nobreak\ W}$ and $p_{\text{blk}}=0.15$.
Figure 6: Tracking performance for the proposed dual-domain framework versus the time index with $\sigma_{z}^{2}=10^{-6}\mathrm{\leavevmode\nobreak\ W}$ and $p_{\text{blk}}=0.15$.
(a) $\sigma_{z}^{2}=10^{-9}\mathrm{\leavevmode\nobreak\ W}$
(b) $\sigma_{z}^{2}=10^{-6}\mathrm{\leavevmode\nobreak\ W}$
Figure 7: CDF of B-Domain multipath AoA tracking errors with $p_{\text{blk}}=0.15$ at different $\sigma_{z}^{2}$ values.
Figure 8: Multipath AoA tracking performance with $\sigma_{z}^{2}=10^{-6}\mathrm{\leavevmode\nobreak\ W}$.
Figure 9: Average AoA tracking errors with respect to the number of antennas.
Figure 10: AoA tracking errors of different predictive BF schemes.
Fig. 5 illustrates the dual-domain tracking performance in terms of root mean squared error (RMSE) for the vehicle position and the absolute error for AoA estimation for path 1. In contrast to the baseline scheme, the proposed dual-domain EKF tracking framework ensures continuous target tracking and achieves a higher tracking accuracy even during the prolonged LoS-absent period. This is because our proposed scheme comprehensively models of the beam transition conditions and effectively utilizes the CKM to extract NLoS path information for positioning. Moreover, it can be observed that all the position tracking RMSE curves first increase and then decrease/stabilize with the evolution of time. This behavior is due to that the tracking accuracy of EKF algorithm requires several iterations to stabilize. Additionally, when path 1 is randomly blocked, the AoA of the strongest NLoS path changes unpredictably leading to beam misalignment and spikes in AoA tracking error indicated by asterisks in Fig. 5. Fortunately, these spikes do not significantly affect overall position tracking accuracy, as the measurement models in the C-Domain for LoS-present and LoS-absent cases are different. Note that the multipath AoA estimation accuracy has not been provided in Fig. 5 since the baseline only tracks the AoA of the LoS path. In the first few time slots, the AoA of LoS path in the following time slots of the baseline scheme is larger than that of the proposed framework because the MAP-based update principle proposed in Section V-C effectively incorporates a priori information when echo signal measurements are unreliable or have low SNR. In contrast, large errors in both target and beam tracking for the baseline scheme arise with time slot, even reaching more than $10$ meters of RMSE of position tracking and $50$ degrees of path 1 AoA tracking error in the end. This is because for LoS-absent cases, the echo signal measurement is treated as meaningless in the baseline scheme in [4] and relies on the state evolution model to predict the position of the vehicle and AoA of LoS path in the following time slots. Since the state evolution model adopted may suffer from model mismatch, with the increase of the slot, both the target and beam tracking error scales up.
Fig. 6 considers a higher noise power to evaluate the performance of the proposed scheme. When the noise power is higher, a larger antenna array can improve tracking accuracy due to a higher BF gain, but they also increase the likelihood of beam misalignment, as shown in Fig. 6. For a larger antenna array with $N_{t}=N_{r}=64$, both target and beam tracking performance becomes worse when the vehicle is close to the RSU, specifically between $80-100$ slots, even for the cases where path 1 has not been blocked. In fact, the AoA of path 1 varies rapidly in these time slots and B-Domain state prediction becomes inaccurate, and thus predictive BF with a narrow beam for a larger antenna array is more susceptible to missing the moving target.
Note that in both Fig. 5 and Fig. 6, baseline methods fail to maintain accuracy tracking in the considered complex and high-mobility wireless network due to random path blockage. The proposed CKM-assisted dual-domain tracking framework outperforms the baseline for two key reasons. First, based on the C-Domain tracking results, the CKM provides additional a priori information for B-Domain tracking, which improves AoA tracking accuracy, even under poor channel conditions. Second, conventional EKF-based target tracking schemes, such as those in prior works [7, 5, 4, 43], are highly sensitive to model mismatch errors due to discrepancy between the assumed measurement model and the actual propagation environment. In contrast, the proposed framework utilizes the CKM to serve as the measurement function in the LoS-absent case in addition to the traditional geometric-based measurement model in the LoS-present case.
Furthermore, Fig. 7a and Fig. 7b illustrate the cumulative distribution function (CDF) of AoA tracking error for path 1 and path 2 for different configurations of $N_{t}$ and $N_{r}$, corresponding to Figs. 5 and 6, respectively. Due to the lower noise power $\sigma_{z}^{2}$, Fig. 7a consistently exhibits better AoA estimation accuracy than Fig. 7b for all paths and antenna configurations. In contrast, Fig. 7b reveals that a higher measurement noise power significantly impacts the AoA tracking accuracy, particularly for path 2. This degradation is most pronounced during non-stationary and unpredictable beam state transitions, where the AoA estimation relies solely on the current observed signal and thus is more sensitive to noise-induced distortions. Consistent with the observations in Fig. 6, increasing the number of antennas is not always beneficial for AoA tracking, particularly for path 2 (NLoS path) and the case with a higher noise power. As the NLoS path has a higher random blockage probability, a smaller antenna array with a wider predictive beam has a higher probability of covering the bursty changed AoA of path 2 and thus may provide a high-quality signal measurement in the B-Domain. It is important to note that $p_{\text{blk}}=0.15$ indicates that $85\%$ of the AoA estimates for path 1 are obtained with the assistance of a priori information and $72.25\%$ of the AoA estimates for path 2 are supported with a priori information. In both Figs. 7a and 7b, the AoA tracking error is lower than $10$ degrees with the assistance of a priori information. This indicates that incorporating a priori information significantly improves AoA tracking accuracy, even under low SNR conditions, especially for the NLoS path. This finding underscores the robustness of our proposed approach in maintaining accurate angle estimates, even in the presence of significant noise and high-speed moving target. Fig. 10 further corroborates the analysis above, demonstrating that while larger antenna arrays can provide a better AoA tracking accuracy in stationary or predictable cases, they are more sensitive to unexpected beam state transition in the propagation environment.
Fig. 10 further demonstrates that nonstationary and unpredictable transition $(III)$ degrades the AoA tracking performance compared to the other two types of beam state transition, particularly for NLoS path and the case with a larger antenna array. In fact, in the third type of beam state transition, without a prior information of channel and relying solely on the received signal, it is challenging to estimate the AoA for the NLoS path due to its much higher propagation loss than that of LoS path. Moreover, in this case, the RSU struggles to predict the beam and predictive BF may suffer from the beam misalignment with a higher probability for a larger antenna array.
VII-C Predictive BF Performance
Fig. 10 illustrates that incorporating predictive BF and optimizing power allocation (PA) significantly improves multipath AoA tracking accuracy. For path 1, Fig. 10 shows that BF with equal PA achieves the highest tracking accuracy, followed by predictive BF with optimized PA, and thirdly, the case without predictive BF. For path 2, predictive BF with optimized PA achieves the highest AoA tracking accuracy, followed by predictive BF with equal PA, and lastly, the case without predictive BF. The optimization of power allocation balances the sensing performance across different propagation paths. By adopting the power distribution among multiple beams, the proposed method effectively improves the AoA tracking performance of the path with a lower path gain. This is particularly important for NLoS paths, where predictive BF with optimized PA allows for enhanced tracking performance by compensating its propagation loss. Overall, the combination of predictive BF and optimized PA offers a more robust solution for multipath AoA tracking in complex environments.
VIII Conclusion
In this paper, we proposed a CKM-assisted dual-domain tracking and predictive BF framework for high-mobility wireless networks. In the C-Domain, we proposed an EKF-based target tracking method that is effective for both LoS-present and LoS-absent conditions. In the B-Domain, we developed an angular TPM that combines the channel temporal correlation and a priori information from CKM for each path to assist AoA tracking. Moreover, we analyzed the CRB of AoA estimation and proposed a predictive BF and power allocation design to minimize the maximum AoA estimation error across multiple paths for the next time slot. Simulation results confirm that CKM significantly enhances both the target and beam tracking accuracy, especially in complex high-mobility wireless communication networks. Additionally, incorporating predictive BF and optimized power allocation further improves AoA tracking performance, especially for NLoS paths with high propagation loss. This framework can be expanded to multi-user and multi-UAV networks, accommodating the simultaneous movement of multiple nodes. The proposed technique is a pivotal advancement for the development of intelligent transportation systems, offering improved communication reliability and efficiency in 6G networks.